In the circuit below V1=(100∠0∘) V and V2=(100∠90∘) V. Find the complex power absorbed by:
(a) V1
(b) V2
(c) 6Ω
(d) j4Ω
(e) −j10Ω
(f) 5Ω

5Ω 6Ω V2 V1 -j1OQ

(a) SV1 =  +j  VA

(b) SV2 =  +j  VA

(c) S6Ω =  +j  VA

(c) Sj4Ω =  +j  VA

(d) S−j10Ω =  +j  VA

(e) S5Ω =  +j  VA

Please do not round. Thanks in advance.
Answer #1

Consider the circuit diagram. v 100Z0° V 10020” V :/ 110 9 100290 v -j10 2 100290° V - +- Apply KCL at node a va - ViVa t Va-

Find current 11- I,= va 6+4j (10020°)–(53.3510242.676°) 6+4j = 4.2308–8.8479; =9.80742-64.44° A Find current I, -10j 53.3510

(a) Find the complex power absorbed by V. 100 9.8074 -5 °2206264.44° = 490.372-115.56° VA =-211.573 – 442.37j VA (b) Find the

(d) Find the complex power absorbed by j4 2. S;4n = Vj4snymstiems = (( j4)Iims ) Isms Y 9.8074 (( j4)(9.8074)-64.440 (9.84 -